OpenStudy (anonymous):

Need a second opinion. Two brother leave from the same point driving in opposite directions, the first brother driving at 70mph and the second at 55mph. The first brother has a one-hour head start. How long will they be able to talk on their car phones if the phones have a 300-mile range.

OpenStudy (anonymous):

Let t represent the time the second brother travelled.\[\color{green}{d_1 = 70 (t+1) \quad d_2 = 50 t \\ d_1 + d_2 = 300\\ 70(t + 1) + 50t = 300\\ 70t + 70 + 50t = 300 \\ t = 1.92 } \] The total time is 2.92 h after the first brother began travelling.

OpenStudy (anonymous):

Thanks, I did the same thing, got the same answer but forgot to add the 1hour into the t to find the final solution.

OpenStudy (vincent-lyon.fr):

Why did you use 50 and not 55?

OpenStudy (anonymous):

OOps! I should have used 55 as @Vincent-Lyon.Fr has pointed out correctly ... \[ \color{green}{d_1 = 70 (t+1) \quad d_2 = 55 t \\ d_1 + d_2 = 300\\ 70(t + 1) + 55t = 300\\ 70t + 70 + 55t = 300 \\ t = 1.84 } \]